In a store, a 40% discount was applied to the price P of an item. In another store, a 50% discount was applied to the price Q of the same item. In both stores the final price after applying the discount was $18. What is the difference, in dollars, between prices P and Q?
Answer by math_tutor2020(3817)


P and Q are the original prices before the 40% and 50% discount respectively.

If you save 40%, then you pay the remaining 60%
0.60P = price after discount
0.60P = 18
P = 18/0.60
P = 30

If you save 50%, then you pay the remaining 50%
0.50Q = 18
Q = 18/0.50
Q = 36

The difference between the original prices is Q-P = 36-30 = 6 dollars